OpenStudy (paystub-2):

how would i graph y=-2x ?

OpenStudy (anonymous):

what do you know about y=-2x?

OpenStudy (anonymous):

slope = ? y -intercept = ?

OpenStudy (paystub-2):

well to tell yhu i already did most of my homework on that its just i need help on 3 more including this one which are y=-x and y=2/3x + 4 so as yu can see all three of them are similar and thats what i need help on.

OpenStudy (anonymous):

x=0; y=0 (0,0) x=1;y=-2 (1,-2) connect those to points

OpenStudy (paystub-2):

slpoe = -2 and intercept = 0 i think

OpenStudy (carson889):

The slope is -2. And with the slope defined as the rise over run, for each push in the positive x-direction the point will go DOWN by 2 in the y-axis. So start at (0,0), then at x = 1, y will equal -2. At x = 2, y will equal -4. For x = -1, y will equal 2. It can be said that for each point along the x-axis the y-value at that point will equal negative two times the x-value.

OpenStudy (anonymous):

yes. you can start the graph at 0 going downwards. slope = y/x so... slope = -2/1 the graph should end up looking like a backward slash on your keyboard.

OpenStudy (paystub-2):

so down 2, right 1?

OpenStudy (anonymous):

yes